# On the Regularization of Learnable Embeddings for Time Series Processing

### v1

@@ Abstract
In processing multiple time series, accounting for the individual features of each sequence can be challenging. To address this, modern deep learning methods for time series analysis combine a shared (global) model with local layers, specific to each time series, often implemented as learnable embeddings. Ideally, these local embeddings should encode meaningful representations of the unique dynamics of each sequence. However, when these are learned end-to-end as parameters of a forecasting model, they may end up acting as mere sequence identifiers. Shared processing blocks may then become reliant on such identifiers, limiting their transferability to new contexts. In this paper, we address this issue by investigating methods to regularize the learning of local learnable embeddings for time series processing. Specifically, we perform the first extensive empirical study on the subject and show how such regularizations consistently improve performance in widely adopted architectures. Furthermore, we show that methods preventing the co-adaptation of local and global parameters are particularly effective in this context. This hypothesis is validated by comparing several methods preventing the downstream models from relying on sequence identifiers, going as far as completely resetting the embeddings during training. The obtained results provide an important contribution to understanding the interplay between learnable local parameters and shared processing layers: a key challenge in modern time series processing models and a step toward developing effective foundation models for time series.
@@ 1 Introduction
Among different methods, research has addressed the problem by looking into hybrid global-local architectures, i.e., architectures that combine global models with local learnable components (e.g., layers) specific to a target time series (Wang et al., 2019; Smyl, 2020). As a prominent example, Smyl (2020) won the M4 competition (Makridakis et al., 2020) by combining a shared recurrent neural network (RNN) with local exponential smoothing models fitted on each time series. In the context of spatiotemporal data, research has shifted towards the adoption of learnable embeddings (Bai et al., 2020; Shao et al., 2022), i.e., vectors of learnable parameters, to reduce the cost of learning (more complex) local processing blocks (Cini et al., 2023b). Each embedding, associated with a specific time series, is fed into the shared modeling architecture and trained end-to-end alongside it. These representations can go beyond simply encoding coordinates, as they can account for the dynamics of each time series w.r.t. the other sequences in the collection (Cini et al., 2023b). Parameters of this kind are analogous to the word embeddings used in natural language processing (NLP) (Mikolov et al., 2013; Peng et al., 2015). We will refer to them as local embeddings.
@@
While local embeddings offer significant advantages, their adoption introduces potential drawbacks. In the first place, differently from NLP applications that operate on fixed size dictionary, the target time series might change over time, and new sequences might be added to the collection. Secondly, since embeddings are learned jointly with the entire architecture, co-adaptation with the shared layers might result in embeddings being used as simple sequence identifiers (Geirhos et al., 2020). This interdependence can hinder the flexibility of the shared processing blocks and limit their transferability. Existing works (Yin et al., 2022; Cini et al., 2023b; Prabowo et al., 2024) show evidence that constraining the structure of the embedding space can lead to transferability improvements. However, no prior work has systematically addressed and evaluated regularization methods for local embeddings in forecasting architectures.
@@
To fill this void, we investigate the impact of regularizing the learning of local embeddings for time series within a selection of commonly used deep learning architectures. In particular, we perform an extensive empirical study comparing a variety of regularization strategies, ranging from standard approaches, such as weight penalties and dropout, to more advanced methods, e.g., clustering and variational regularization. Our analysis considers a range of scenarios, including transductive settings, transfer learning, and sensitivity analyses. Empirical results confirm that methods that prevent the co-adaptation of the local and global blocks are consistently among the best-performing approaches. To further validate this observation, we include in the analysis a forget-and-relearn strategy (Zhou et al., 2021) that periodically resets embeddings during training. Results w.r.t. such baseline further show that decoupling the learning of local and global parameters offers a good design principle for regularization strategies in this context. Our main findings can be summarized as follows.
@@
    In both transductive and transfer learning settings, regularizing the local embeddings can provide clear performance improvements across several architectures.
@@
    Regularization strategies that prevent the co-adaptation between global and local parameters consistently lead to larger performance gains.
@@
    Finding (F2) can be used as a design principle for designing new regularization strategies to mitigate overfitting and improve transferability.
The regularization of local learnable embeddings emerges as an often neglected but central aspect, which, with negligible computational overhead, can lead to significant performance improvements across all the considered benchmarks. This makes our study an important missing piece for guiding practitioners and researchers in building and designing modern neural architectures for time series processing.
@@ 2.1 Problem settings
Consider a collection ${\mathcal{D}}$ of $N$ time series, where each $i$-th time series consists of a sequence of $T$ observations $\{{\bm{x}}^{i}_{t}\in{\mathbb{R}}^{d_{x}}\}_{t=1}^{T}$. Specifically, we indicate as related time series a set of homogenous time series coming from the same domain but generated by different sources (e.g., different sensors). Examples include sales figures for different products, or energy consumption of various users. Time series might be acquired both asynchronously and/or synchronously; in the latter case, we denote the stacked $N$ observations at time step $t$ by the matrix ${\bm{X}}_{t}\in{\mathbb{R}}^{N\times d_{x}}$ and use the shorthands ${\bm{X}}_{t:t+T}$ to indicate the sequence of observations within the time interval $[t,t+T)$ and ${\bm{X}}_{<t}$ to indicate those up to time $t$, excluding $t$. Eventual (exogenous) covariates associated with each time series are denoted by ${\bm{u}}^{i}_{t}\in{\mathbb{R}}^{d_{u}}$ (${\bm{U}}_{t}\in{\mathbb{R}}^{N\times d_{u}}$). We focus on multistep-ahead time series forecasting, i.e., the problem of predicting the next $H$ future values for each $i$-th time series ${\bm{x}}^{i}_{t:t+H}$ given exogenous variables and a window of $W$ past observations. We focus on point forecasts, while probabilistic predictors might also be considered.
@@ 2.2 Hybrid global-local architectures for time series
$\hat{\bm{x}}^{i}_{t:t+H}=F({\bm{x}}^{i}_{t-W:t},{\bm{u}}^{i}_{t-W:t+H};\theta),\quad i=1,...,N$ (1)
where $\theta$ are the learnable parameters of the model, $\hat{\bm{x}}^{i}_{t:t+H}$ indicates the predicted values of ${\bm{x}}^{i}_{t:t+H}$ and $F(\cdot;\theta)$ is a family of parametric forecasting models. Models in Eq. 1 do not take into account (spatial) dependencies that might exist among time series. In scenarios such as spatiotemporal forecasting, where such dependencies might be relevant to achieve accurate predictions, models that can operate on sets of (synchronous) correlated time series should be preferred. In particular, we consider a family of models
$\widehat{\bm{X}}^{{\mathcal{S}}}_{t:t+H}=F({\bm{X}}^{{\mathcal{S}}}_{t-W:t},\mathbf{U}^{{\mathcal{S}}}_{t-W:t+H};\theta),\quad\quad\forall{\mathcal{S}}\subseteq{\mathcal{D}}$ (2)
where $F(\cdot;\theta)$ operates on subsets ${\mathcal{S}}$ of the collection of time series ${\mathcal{D}}$ and ${\bm{X}}_{t}^{{\mathcal{S}}}\in{\mathbb{R}}^{|{\mathcal{S}}|\times d_{x}}$ the resulting stack of observations at time step $t$. Models of this kind can be implemented by architectures operating on sets (Zaheer et al., 2017), e.g., attention-based architectures (Vaswani et al., 2017; Grigsby et al., 2021) or spatiotemporal graph neural networks (STGNNs) (Jin et al., 2023; Cini et al., 2023a) based on message passing (Gilmer et al., 2017). These models can eventually account for priors on existing dependencies among time series, which could be encoded, e.g., by a graph adjacency matrix $\mathbf{A}\in{\mathbb{R}}^{N\times N}$ (Cini et al., 2023a). Note that models in both families (Eqs. 1 and 2) are global, i.e., they share parameters among all the time series being processed, with clear advantages in terms of sample efficiency (Montero-Manso & Hyndman, 2021). Most of the recent successes in applying deep learning to time series forecasting are based on the idea of implementing such global models with a neural network (Benidis et al., 2022). With full generality, in the following, we will consider models belonging to the family in Eq. 2, as Eq. 1 can be seen as a special case where $|{\mathcal{S}}|=1$. In particular, we consider the whole collection at once (i.e., $|{\mathcal{S}}|=N$) and drop the superscript ${\mathcal{S}}$ to simplify the notation.
@@
${\bm{H}}_{t}^{l+1}=\textsc{PropLayer}^{l}\left({\bm{H}}_{<t}^{l}\right),\quad l=0,1,...,L-1$ (4)
@@
where the additional index $l$ refers to the layer depth, ${\bm{h}}^{i,l}_{t}\in{\mathbb{R}}^{d_{h}}$ is the representation associated with the $i$-th time series at the $l$-th layer, and ${\bm{H}}^{l}_{t}\in{\mathbb{R}}^{|{\mathcal{S}}|\times d_{h}}$ the stack of such representations. The $\textsc{Encoder}\left(\cdot\right)$ and $\textsc{Decoder}\left(\cdot\right)$ blocks can be implemented in different ways (e.g. a linear layer or a multilayer perceptron (MLP)) and operate w.r.t. single time steps and time series. Conversely, the $\textsc{PropLayer}\left(\cdot\right)$ blocks are the only compontens of the architecture propagating representations across the temporal dimension and across time series in the collection. Each $\textsc{PropLayer}\left(\cdot\right)$ can be implemented by any existing sequence modeling architecture, e.g., RNNs (Hochreiter & Schmidhuber, 1997; Cho et al., 2014a) or temporal convolutional networks (TCNs) (LeCun & Bengio, 1998; Bai et al., 2018), and/or spatiotemporal operators, e.g., spatiotemporal attention (STAtt) models (Grigsby et al., 2021; Deihim et al., 2023). In particular, STGNNs (Seo et al., 2018; Yu et al., 2018), i.e., models that combine sequence modeling architectures with message passing operators, are among the most popular architectures for implementing propagation layers when processing correlated time series (Jin et al., 2023).
@@
As anticipated in the introduction, despite their advantages, global models might struggle to account for the specific dynamics of each time series. Hybrid global-local architectures (Benidis et al., 2022) address this issue by including parameters specific to each target time series. If we indicate these parameters as $\Phi=\{\phi^{0},...,\phi^{N}\}$, the resulting model family would provide forecasts for the time series in the collection as
$\widehat{\bm{X}}^{\mathcal{S}}_{t:t+H}=F\left({\bm{X}}^{\mathcal{S}}_{t-W:t},{\bm{U}}^{\mathcal{S}}_{t-W:t+H};\theta,\Phi^{\mathcal{S}}\right).$ (6)
Although many implementations of such models exist (Benidis et al., 2022; Cini et al., 2023b), we consider models where the local components are implemented as embeddings ${\bm{E}}\in{\mathbb{R}}^{N\times d_{e}}$ of learnable parameters such that $\Phi={\bm{E}}$ and $\phi^{i}={\bm{e}}^{i}$. In particular, each local embedding $\mathbf{e}^{i}$ is associated with the corresponding $i$-th time series and can be learned end-to-end jointly with the shared network weights. Embeddings are incorporated into the template architecture at both the encoder and decoder level by concatenating them to the input as
@@ 3 Related works
Learnable embeddings are key components of state-of-the-art time series processing architectures such as STGNNs (Bai et al., 2020; Cini et al., 2023a) and attention-based models (Marisca et al., 2022; Liu et al., 2023; Xiao et al., 2024). In particular, Cini et al. (2023b) systematically addresses the role of such embeddings in hybrid global-local STGNNs. Aside from modeling local dynamics, embeddings have been used extensively to amortize the cost of learning the full adjacency matrix in graph-based models (Wu et al., 2019; Shang et al., 2021; Satorras et al., 2022; De Felice et al., 2024), or as spatial positional encodings (Marisca et al., 2022; Shao et al., 2022; Liu et al., 2023). They are also routinely used in NLP as word embeddings (Mikolov et al., 2013) and, often, as spatial positional encodings (Devlin et al., 2019; Yang et al., 2019).
@@ 4 Regularization strategies for local embeddings
In this section, we discuss possible shortcomings of hybrid global-local forecasting architectures and present the regularization strategies employed in our experimental analysis.
Figure 2: Validation curves when regularizing whole model or just local parameters in a time series forecasting task (50 runs, ±$\pm 1\sigma$).
Global models are inherently less likely to overspecialize to individual sequences as, having limited capacity, it would compromise performance w.r.t. other sequences within the set. The introduction of local embeddings reduces the regularization effect brought by this inductive bias, resulting in a model that may become more susceptible to overfitting to individual time series, with a potential negative impact on both performance and transferability (Cini et al., 2023b). Model regularization is a common approach to deal with overfitting and to control model and sample complexity. For instance, one might apply well-known techniques, such as weight decay (Krogh & Hertz, 1991) or dropout (Srivastava et al., 2014), across the entire network. However, in many scenarios, the global processing blocks are not particularly overparametrized, as weights are shared. Conversely, in global-local models, one could consider regularizing the local parameters only, i.e., in our settings, constraining how embeddings are learned. Fig. 2 empirically supports this intuition with an example in a simple setting: given a global time-series forecasting model with fixed hyperparameters (green), adding local embeddings (orange) improves performance, but makes it more likely to overfit (as model complexity increases). Regularizing the entire global-local model (purple), however, hinders performance in this scenario. Conversely, applying regularization (dropout in this case) to the embeddings only (red), regularizes the model effectively. Note that regularization of the global parameters can obviously still be beneficial in the case of an over-parametrized global model.

### v2

# On the Regularization of Learnable Embeddings for Time Series Forecasting

@@ Abstract
In forecasting multiple time series, accounting for the individual features of each sequence can be challenging. To address this, modern deep learning methods for time series analysis combine a shared (global) model with local layers, specific to each time series, often implemented as learnable embeddings. Ideally, these local embeddings should encode meaningful representations of the unique dynamics of each sequence. However, when these are learned end-to-end as parameters of a forecasting model, they may end up acting as mere sequence identifiers. Shared processing blocks may then become reliant on such identifiers, limiting their transferability to new contexts. In this paper, we address this issue by investigating methods to regularize the learning of local learnable embeddings for time series processing. Specifically, we perform the first extensive empirical study on the subject and show how such regularizations consistently improve performance in widely adopted architectures. Furthermore, we show that methods attempting to prevent the co-adaptation of local and global parameters by means of embeddings perturbation are particularly effective in this context. In this regard, we include in the comparison several perturbation-based regularization methods, going as far as periodically resetting the embeddings during training. The obtained results provide an important contribution to understanding the interplay between learnable local parameters and shared processing layers: a key challenge in modern time series processing models and a step toward developing effective foundation models for time series.
@@ 1 Introduction
Among different methods, research has addressed the problem by looking into hybrid global-local architectures, i.e., architectures that combine global models with local learnable components (e.g., layers) specific to a target time series (Wang et al., 2019; Smyl, 2020). As a prominent example, Smyl (2020) won the M4 competition (Makridakis et al., 2020) by combining a shared recurrent neural network (RNN) with local exponential smoothing models fitted on each time series. In the context of spatiotemporal data, research has shifted towards the adoption of learnable embeddings (Bai et al., 2020; Shao et al., 2022), i.e., vectors of learnable parameters, to reduce the cost of learning (more complex) local processing blocks (Cini et al., 2023b). Each embedding, associated with a specific time series, is fed into the shared modeling architecture and trained end-to-end alongside it. These representations can go beyond simply encoding coordinates, as they can account for the dynamics of each time series with respect to the other sequences in the collection (Cini et al., 2023b). Parameters of this kind are analogous to the word embeddings used in natural language processing (NLP) (Mikolov et al., 2013; Peng et al., 2015). We will refer to them as local embeddings.
@@
While local embeddings offer significant advantages, their adoption introduces potential drawbacks. In the first place, differently from NLP applications that operate on a fixed-size dictionary, the target time series might change over time, and new sequences might be added to the collection. Secondly, since embeddings are learned jointly with the entire architecture, co-adaptation (Srivastava et al., 2014) with the shared layers might result in embeddings being used as simple sequence identifiers (Geirhos et al., 2020). This interdependence can hinder the flexibility of the shared processing blocks and limit their transferability. Existing works (Yin et al., 2022; Cini et al., 2023b; Prabowo et al., 2024) show evidence that constraining the structure of the embedding space can lead to transferability improvements. However, no prior work has systematically addressed and evaluated regularization methods for local embeddings in forecasting architectures.
@@
To fill this void, we investigate the impact of regularizing the learning of local embeddings for related time series within a selection of commonly used deep learning forecasting architectures. In particular, we perform an extensive empirical study comparing a variety of regularization strategies, ranging from standard approaches, such as weight penalties and dropout, to more advanced methods, e.g., clustering and variational regularization. Our analysis considers a range of scenarios, including transductive settings, transfer learning, and sensitivity analyses. Empirical results show that methods that attempt to prevent the co-adaptation of the local and global blocks by perturbing the embeddings at training time are consistently among the best-performing approaches. To further validate this observation, we include in the analysis a forget-and-relearn strategy (Zhou et al., 2021), named forgetting regularization, that periodically resets embeddings during training. The results obtained by considering this approach further show that perturbation of local parameters offers a good design principle for regularization strategies in this context. Our main findings can be summarized as follows.
@@
    In both transductive and transfer learning settings, regularizing the local embeddings can provide consistent performance improvements across several forecasting architectures.
@@
    Regularization strategies based on local parameters’ perturbations aimed at preventing co-adaptation consistently lead to larger performance gains compared to other approaches.
@@
    Finding (F2) can be used as a design principle for designing new regularization strategies to mitigate overfitting and improve transferability, as highlighted by the competitive performance of the considered forgetting regularization.
The regularization of local learnable embeddings emerges as an often neglected but central aspect, which, with negligible computational overhead, can lead to performance improvements across all the considered benchmarks. This makes our study an important missing piece for guiding practitioners and researchers in building and designing modern neural architectures for time series processing.
@@ 2.1 Problem settings
Consider a collection ${\mathcal{D}}$ of $N$ time series, where each $i$-th time series consists of a sequence of $T$ observations $\{{\bm{x}}^{i}_{t}\in{\mathbb{R}}^{d_{x}}\}_{t=1}^{T}$. Specifically, we indicate as related time series a set of homogenous time series coming from the same domain but generated by different sources (e.g., different sensors). Examples include sales figures for different products, or energy consumption of various users. Time series might be acquired both asynchronously and/or synchronously; in the latter case, we denote the stacked $N$ observations at time step $t$ by the matrix ${\bm{X}}_{t}\in{\mathbb{R}}^{N\times d_{x}}$. We use the shorthands ${\bm{X}}_{t:t+T}$ to indicate the sequence of observations within the time interval $[t,t+T)$ and ${\bm{X}}_{\leq t}$ to indicate those up to time $t$ (included). Eventual (exogenous) covariates associated with each time series are denoted by ${\bm{u}}^{i}_{t}\in{\mathbb{R}}^{d_{u}}$ (${\bm{U}}_{t}\in{\mathbb{R}}^{N\times d_{u}}$). We focus on multistep-ahead time series forecasting, i.e., the problem of predicting the next $H$ future values for each $i$-th time series ${\bm{x}}^{i}_{t:t+H}$ given exogenous variables and a window of $W$ past observations. We focus on point forecasts, while probabilistic predictors might also be considered.
@@ 2.2 Hybrid global-local architectures for time series
$\hat{\bm{x}}^{i}_{t:t+H}=F({\bm{x}}^{i}_{t-W:t},{\bm{u}}^{i}_{t-W:t+H};\,\bm{\theta}),\quad i=1,...,N$ (1)
where $\bm{\theta}$ are the learnable parameters of the model, $\hat{\bm{x}}^{i}_{t:t+H}$ indicates the predicted values of ${\bm{x}}^{i}_{t:t+H}$, and $F(\cdot;\bm{\theta})$ is a family of parametric forecasting models. Models in Eq. 1 do not take into account (spatial) dependencies that might exist among time series. In scenarios such as spatiotemporal forecasting, where such dependencies might be relevant to achieve accurate predictions, models that can operate on sets of (synchronous) correlated time series should be preferred. In particular, we consider a family of models
$\widehat{\bm{X}}^{{\mathcal{S}}}_{t:t+H}=F({\bm{X}}^{{\mathcal{S}}}_{t-W:t},\mathbf{U}^{{\mathcal{S}}}_{t-W:t+H};\,\bm{\theta}),\quad\quad\forall{\mathcal{S}}\subseteq{\mathcal{D}}$ (2)
where $F(\cdot;\bm{\theta})$ operates on subsets ${\mathcal{S}}$ of the collection of time series ${\mathcal{D}}$, and ${\bm{X}}_{t}^{{\mathcal{S}}}\in{\mathbb{R}}^{|{\mathcal{S}}|\times d_{x}}$ is the corresponding stack of observations at time step $t$. Models of this kind can be implemented by architectures operating on sets (Zaheer et al., 2017), e.g., attention-based architectures (Vaswani et al., 2017; Grigsby et al., 2021) or spatiotemporal graph neural networks (STGNNs) (Jin et al., 2023; Cini et al., 2023a) based on message passing (Gilmer et al., 2017). These models can eventually account for priors on existing dependencies among time series, which could be encoded, e.g., by a graph adjacency matrix $\mathbf{A}\in{\mathbb{R}}^{N\times N}$ (Cini et al., 2023a). Note that models in both families (Eqs. 1 and 2) are global, i.e., they share parameters among all the time series being processed, with clear advantages in terms of sample efficiency (Montero-Manso & Hyndman, 2021). Most of the recent successes in applying deep learning to time series forecasting are based on the idea of implementing such global models with a neural network (Benidis et al., 2022). In the following, we will consider models belonging to the family in Eq. 2, as Eq. 1 can be seen as a special case where $|{\mathcal{S}}|=1$. In particular, we consider the whole collection at once (i.e., $|{\mathcal{S}}|=N$) and drop the superscript ${\mathcal{S}}$ to simplify the notation.
@@
${\bm{H}}_{t}^{l+1}=\textsc{PropLayer}^{l}\left({\bm{H}}_{\leq t}^{l}\right),\quad l=0,1,...,L-1$ (4)
@@
where the additional index $l$ refers to the layer depth, ${\bm{h}}^{i,l}_{t}\in{\mathbb{R}}^{d_{h}}$ is the representation associated with the $i$-th time series at the $l$-th layer, and ${\bm{H}}^{l}_{t}\in{\mathbb{R}}^{|{\mathcal{S}}|\times d_{h}}$ the stack of such representations. The $\textsc{Encoder}\left(\cdot\right)$ and $\textsc{Decoder}\left(\cdot\right)$ blocks can be implemented in different ways, e.g. a linear layer or a multilayer perceptron (MLP), and operate on single time steps and time series. Conversely, the $\textsc{PropLayer}\left(\cdot\right)$ blocks are the only components of the architecture propagating representations across the temporal dimension and across time series in the collection. Each $\textsc{PropLayer}\left(\cdot\right)$ can be implemented by any existing sequence modeling architecture, e.g., RNNs (Hochreiter & Schmidhuber, 1997; Cho et al., 2014a) or temporal convolutional networks (TCNs) (LeCun & Bengio, 1998; Bai et al., 2018), and/or spatiotemporal operators, e.g., spatiotemporal attention (STAtt) models (Grigsby et al., 2021; Deihim et al., 2023). In particular, STGNNs (Seo et al., 2018; Yu et al., 2018), i.e., models that combine sequence modeling architectures with message passing operators, are among the most popular architectures for implementing propagation layers when processing correlated time series (Jin et al., 2023).
@@
As anticipated in the introduction, despite their advantages, global models might struggle to account for the specific dynamics of each time series. Hybrid global-local architectures (Benidis et al., 2022) address this issue by including parameters specific to each target time series. If we indicate these parameters as $\Phi=\{\bm{\phi}^{1},...,\bm{\phi}^{N}\}$, the resulting model family would provide forecasts for the time series in the collection as
$\widehat{\bm{X}}^{\mathcal{S}}_{t:t+H}=F\left({\bm{X}}^{\mathcal{S}}_{t-W:t},{\bm{U}}^{\mathcal{S}}_{t-W:t+H};\,\bm{\theta},\Phi^{\mathcal{S}}\right).$ (6)
Although many implementations of such models exist (Benidis et al., 2022; Cini et al., 2023b), we consider models where the local components are implemented as embeddings ${\bm{E}}\in{\mathbb{R}}^{N\times d_{e}}$ of learnable parameters such that $\Phi={\bm{E}}$ and $\bm{\phi}^{i}={\bm{e}}^{i}$. In particular, each local embedding $\mathbf{e}^{i}$ is associated with the corresponding $i$-th time series and can be learned end-to-end jointly with the shared network weights. Embeddings are incorporated into the template architecture at both the encoder and decoder level by concatenating them to the input as
@@ 3 Related works
Learnable embeddings are key components of state-of-the-art time series processing architectures such as STGNNs (Bai et al., 2020; Cini et al., 2023a) and attention-based models (Marisca et al., 2022; Liu et al., 2023; Xiao et al., 2024). In particular, Cini et al. (2023b) systematically addresses the role of such embeddings in hybrid global-local STGNNs. Aside from modeling local dynamics, embeddings have been used extensively to amortize the cost of learning the full adjacency matrix in graph-based models (Wu et al., 2019; Shang et al., 2021; Satorras et al., 2022; De Felice et al., 2024), or as spatial positional encodings (Marisca et al., 2022; Shao et al., 2022; Liu et al., 2023). They are also routinely used in NLP as word embeddings (Mikolov et al., 2013) and often as spatial positional encodings (Devlin et al., 2019; Yang et al., 2019).
@@ 4 Regularization strategies for local embeddings
In this section, we discuss some possible shortcomings of hybrid global-local forecasting architectures and present the regularization strategies employed in our experimental analysis.
Figure 2: Validation curves when regularizing whole model or just local parameters in a time series forecasting task (STGNN model, METR-LA dataset, 50 runs, ±$\pm 1std$). Models and datasets are discussed in Sec. 5.
Global models are inherently less likely to overspecialize to individual sequences since all parameters are shared among time series. The introduction of local embeddings reduces the regularization effect brought by this inductive bias and jointly learning local and global parameters end-to-end on a downstream task may result in a model that is more likely to overfit to individual time series, with a potential negative impact on both performance and transferability (Cini et al., 2023b). In the following, we use the term co-adaptation to refer to these overfitting phenomena coming from training the local embeddings end-to-end with the global model.
Model regularization is a common approach to deal with overfitting and to control model and sample complexity. For instance, one might apply well-known techniques, such as weight decay (Krogh & Hertz, 1991) or dropout (Srivastava et al., 2014), across the entire network. However, in many scenarios, the global processing blocks are not particularly overparametrized, as weights are shared, and regularization might not be necessary. In global-local models, one could consider regularizing the local parameters only, i.e., in our settings, constraining how embeddings are learned. Fig. 2 empirically supports this intuition with an illustrative example: given a global time-series forecasting model with fixed hyperparameters (green), adding local embeddings (orange) improves performance, but makes it more likely to overfit (as model complexity increases). Regularizing the entire global-local model (purple), however, hinders performance in this scenario. Conversely, applying regularization (dropout in this case) to the embeddings only (red), regularizes the model effectively. Note that regularization of the global parameters can obviously still be beneficial in the case of an over-parametrized global model.